OVHCloud Network Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#14205 — nwk-5-a9
Incident Report for Network & Infrastructure
Resolved
There are OSPF flaps on this harware (BDF session down).

Update(s):

Date: 2015-08-13 12:37:54 UTC
The router is up. We will monitor the BFD flaps to make sure the problem is fixed.

Date: 2015-08-13 12:37:30 UTC
Reboot underway.

Date: 2015-08-13 12:37:22 UTC
Starting to install the router.

Date: 2015-08-12 17:31:30 UTC
We must install an EMS (patch), installation requires a reboot of the equipment.

The operation will take place on August 13th at 10am CEST.

Date: 2015-08-12 17:30:22 UTC
A case was opened at the manufacturer´s in order to investigate.

Date: 2015-08-03 08:39:40 UTC
We can't find any issue with the links: optical links are fine, there aren't any CRC errors.

We will continue to look for the issue.
Posted Aug 03, 2015 - 08:38 UTC